OpenStudy (anonymous):

Number of 4 different digits is formed by using 1, 2, 3, 4, 5, 6, 7. Find the probability that it is divisible by 5.

terenzreignz (terenzreignz):

You can repeat digits, I assume?

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

Well, then, this is just like drawing from a pool containing all these digits, only everytime you draw, you put it back, right?

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

It's like there's an urn, with 7 chips, one marked 1, one marked 2, etc. And you pick a chip from that urn, four times, but every time you pick a chip, you put it back.

terenzreignz (terenzreignz):

So actually, it doesn't matter what the first three digits are, just that the last (fourth) digit must be 5, right?

OpenStudy (anonymous):

is the solution should be liek this: total four digit numbers that can be formed by 1 2 3 4 5 6 7 are 7*6*5*4=840 total four digit numbers that can be formed by 1 2 3 4 5 6 7 where last number is 5 are7*6*5=210 probablity is 210/840=1/4

terenzreignz (terenzreignz):

Question... how did you get that the probability of a four digit number with 5 as a last digit is 5?

OpenStudy (anonymous):

plse tell me how to solve

terenzreignz (terenzreignz):

First of all, how many possible four digit numbers can you form? 8 x 8 x 8 x 8, right, since digits can be repeated?

OpenStudy (anonymous):

k

terenzreignz (terenzreignz):

Nope, I was wrong :D It's just 7 x 7 x 7 x 7 My bad :D

OpenStudy (anonymous):

k

terenzreignz (terenzreignz):

Now, how many of these end with 5? They're 7 x 7 x 7 x 1, right, since you can only have 5 for your last digit.

OpenStudy (anonymous):

k

terenzreignz (terenzreignz):

well, divide \[\large \frac{7\cdot7\cdot7\cdot1}{7\cdot7\cdot7\cdot7}\] What do you get?

OpenStudy (anonymous):



terenzreignz (terenzreignz):

correct